Suppose you have two options when investing money in the stock market: stock A and stock B. The returns on both are dependent on the state of the economy, which fluctuates with the business cycle. During periods of strong economic growth, the rates of return for stock A and stock B are 26.00 and 9.00, respectively. Periods of weak growth during recessions cause the rates of return for stock A and stock B to fall to 3.00 and 1.00, respectively. Additionally, assume that an economic boom is twice as likely as an economic downturn.

Calculate the expected return for stock A:    (Round to two decimals, if necessary.)


Calculate the expected return for stock B:  (Round to two decimals, if necessary.)


1
Expert's answer
2021-10-18T11:28:16-0400

A.

"26-3=23"

"23\times2=46" %

Expected return for stock:

="\frac{26\times146}{100}=37.96"


B.

"9-1=8"

"8\times2=16" %

An increase of 16% implies, 100+16=116%.

Expected return for stock:

"=\frac{9\times116}{100}=10.44"